C Programming Question 13 4 pts 4 pts What is the output of the following code? int x-5; x*= 3; printf(%d\n,x); x=x/2 printf(%d\n,x); printf(%d\n,x++); printf(%d\n,++x);
Answer #1

Initially x is 5

x*=3 that means x = x*3. x contains 5*3 = 15 now.
x++. x is 16
++x, x is 17
x = x/2, x is 8
printf("%d\n",x); will produce 8
printf("%d\n",x++); will produce 8, now x is 9
printf("%d\n",++x); will produce 10, now x is 10

OUTPUT SCREEN IS :